OpenStudy (anonymous):

Please Help Me Solve this! A tank of medical oxygen hold 5.0 litters of O2 at the pressure of 130 atm and a temperature of 25 degrees Celsius. What volume would the same amount of oxygen occupy at 1.0 atm and 25 degrees Celsius?

OpenStudy (anonymous):

This problem is a ideal gas law problem (at least that's how I see it) in chemistry. What we must use is the ideal gas law. Which is PV=nRT where P is pressure in atm, V is volume in L, n is number of moles, R is an arbitrary constant which should be .0821, and T is Temperature in K Given the first little bit of information we receive, we can deduce the correct number of moles. So using PV=nRT we get: ((130 atm)(5 L))/((.0821)(298.15 K)=n(# of moles) The temperature I listed is 298.15 K because Kelvin is the amount of degrees Celsius added to 273.15 (Because -273.15 degrees C is absolute zero and 0 Kelvin is also absolute zero) When we plug this into a calculator, we get the number of moles as 26.55 moles (About) Now in the second part of the question we have to solve for the Volume the oxygen would occupy when we have the same number of moles. So manipulating the ideal gas law again to solve for Volume gives us V=nRT/P Plugging in our new information gives us, V=((26.55 moles)(.0821)(298.15 K))/(1 atm) V=650 Liters I hope this helps! And if you have a conflicting answer then please tell me so I can figure it out, but I'm 99% confident this is correct!

OpenStudy (anonymous):

Thank you so much!